# On the Robustness of Kernel Goodness-of-Fit Tests

### v4

@@ A.1.3 Proof of Proposition˜9
We first show that the moment conditions (i)-(ii) in Lemma˜7 and Lemma˜8 are met, and apply these lemmas to conclude the proof.
Step 1: Verifying moment conditions
For any $\mathbf{x},\mathbf{z}\in\mathbb{R}^{d}$, we can bound the squared Stein kernel as
@@
Step 2: Proving the main claim
Pick any $s\in[0,1)$ and any integer $m^{\prime}\in\big{[}\frac{1}{2}n^{1-s},n\big{)}$. Define $\epsilon_{m,n}\coloneqq m^{\prime}/n$ and $m=n-m^{\prime}$. Let $C,C_{P,1},C_{P,2}$ be the constants defined in Lemma˜8, and define the events
@@ A.1.4 Proof of Theorem˜1
It suffices to show the claim for any sequence of the form $\epsilon_{n}=n^{-s}$, where $s\in[0,1)$ is arbitrary. Let the random variable $M^{\prime}$ be defined as at the start of Section˜A.1. For any random sample $\mathbb{X}_{n}$, define the event $\mathcal{A}(\mathbb{X}_{n})=\{D^{2}(\mathbb{X}_{n})>q^{2}_{\infty,1-\alpha}(\mathbb{X}_{n})\}$. For $n$ sufficiently large so that $n^{(1-s)/2}>2\sqrt{2}C\log(1/\alpha)$, we can apply Proposition˜9 to conclude that, for all $n$, there exists $\mathbf{z}_{n}\in\mathbb{R}^{d}$ such that the following holds for all $m^{\prime}\in\big{[}\frac{1}{2}n^{1-s},n\big{)}=\big{[}\frac{1}{2}\epsilon_{n}n,n\big{)}$,
@@
which will imply the claimed result of our theorem. Using a concentration inequality for Binomial distributions (e.g., Chung and Lu,, 2002, Lemma 2.1) applied to the case of a $\text{Binomial}(n,\epsilon_{n})$, we can show that the event $\mathcal{B}\coloneqq\{\epsilon_{n}n/2\leq M^{\prime}\leq 3\epsilon_{n}n/2\}$ holds with high probability:
@@
$\stackrel{{\scriptstyle(b)}}{{=}}\;\min_{m^{\prime}\in I_{n}}\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}\big{(}\mathcal{A}(\mathbb{X}_{n})\;|\;M^{\prime}=m^{\prime}\big{)}\Pr_{M^{\prime}}(\mathcal{B})$
@@
where the first inequality follows from the law of total probability, $(a)$ and $(b)$ hold by the partition $\mathcal{B}=\cup_{m^{\prime}\in I_{n}}\{M^{\prime}=m^{\prime}\}$ and again the law of total probability, $(c)$ follows from (24) and (26), and $(d)$ holds since the assumptions $\epsilon_{n}=n^{-s}$ and $s\in[0,1)$ imply $f(n)=o\big{(}n^{-(1-s)}\big{)}$. On the other hand, the assumed moment conditions imply $\mathbb{E}_{\mathbf{X}^{\ast}\sim P}[u_{p}(\mathbf{X}^{\ast},\cdot)]=0$ as argued in the paragraph before (11) and $\mathbb{E}_{\mathbf{X}^{\ast}_{1},\mathbf{X}^{\ast}_{2}\sim P}[u_{p}(\mathbf{X}^{\ast}_{1},\mathbf{X}^{\ast}_{2})^{2}]\leq\mathbb{E}_{\mathbf{X}^{\ast}\sim P}[u_{p}(\mathbf{X}^{\ast},\mathbf{X}^{\ast})^{2}]<\infty$ as argued in Section˜A.1.3, so we can apply Arcones and Gine, (1992, Theorem 3.5) with $r=2$ to yield the asymptotic validity of the KSD test when $Q=P$, namely
@@ A.1.5 Implications for KSDs on Non-Euclidean Spaces
The core argument underpinning the proofs in Theorem˜1 is that the function $\mathbf{z}\mapsto u_{p}(\mathbf{z},\mathbf{z})$ is unbounded on the support of the model, which in our case is $\mathbb{R}^{d}$. The proof per se does not rely on the support being an Euclidean space. Consequently, this argument can potentially be extended to KSDs constructed for models defined on other spaces, such as models for Riemannian data (Xu and Matsuda,, 2021), sequence data with varying dimensionality (Baum et al.,, 2023), functional data (Wynne et al.,, 2022), or censored time-to-event data (Fernandez et al.,, 2020).
@@ A.2 Proof of Lemma˜2
Direct expansion of the Stein kernel in (1) with a tilted kernel $k$ gives
@@
where the last line follows from Cauchy-Schwarz inequality. The RHS of (29) is bounded over $\mathbf{x}$ assuming the stated conditions on $h$, $w$ and the supremum of $\|s_{p}(x)w(x)\|_{2}=\|s_{p,w}(x)\|_{2}$. To prove the bound on $u_{p}(\mathbf{x},\mathbf{x}^{\prime})$, we denote by $\mathcal{H}_{u}$ the RKHS associated with $u_{p}$, which is a reproducing kernel (see, for example, Barp et al.,, 2024, Theorem 1). We use the reproducing property of the Stein reproducing kernel $u_{p}$ and the Cauchy-Schwarz inequality for the corresponding RKHS norm $\|\cdot\|_{\mathcal{H}_{u}}$ to yield
@@ A.3 Proof of Theorem˜3 and Related Preliminary Results
  • Section˜A.3.2 shows Lemma˜12, which says that the difference in the exceedance probabilities of $\Delta_{\theta}(\mathbb{X}_{n})$ and of $\Delta_{\theta}(\mathbb{X}_{n}^{\ast})$ is small.
  • Section˜A.3.3 shows Lemma˜13, which says that the rejection probability of a robust-KSD test using $\mathbb{X}_{n}$ is close to that of a test using $\mathbb{X}_{n}^{\ast}$, assuming $\mathbb{X}_{n}$ and $\mathbb{X}_{n}^{\ast}$ differ by no more than $o(n^{1/2})$ elements. Its proof relies on Lemma˜11 and Lemma˜12.
  • Section˜A.3.4 uses Lemma˜13 to show Proposition˜14, which says that the robust-KSD test is qualitatively robust for any $\theta\geq 0$. This result immediately shows Theorem˜3, the proof of which is also contained in this subsection.
@@
Assume $k$ is a tilted kernel satisfying the conditions in Lemma˜2. Then there exists absolute constants $C_{1},C_{2}>0$ such that, for any $\delta>0$ and any (deterministic) realisations $\mathbb{X}_{n}=\{\mathbf{x}_{i}\}_{i=1}^{n}$ and $\mathbb{X}_{n}^{\ast}=\{\mathbf{x}_{i}^{\ast}\}_{i=1}^{n}$ that differ by at most $m^{\prime}$ elements, we have
@@
$\omega(m^{\prime})\;$ $\coloneqq\;\big{|}\Pr\nolimits_{\mathbb{X}_{m}^{\ast}\sim P,\mathbf{W}}\left(\Delta_{\theta}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})>q_{\infty,1-\alpha}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})\right)$
$\;\quad\quad-\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\left(\Delta_{\theta}(\mathbb{X}_{n}^{\ast})>q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})\right)\big{|}\;.$
@@ A.3.1 Proof of Lemma˜11
Without loss of generality, we assume $\mathbb{X}_{n}$ and $\mathbb{X}_{n}^{\ast}$ are ordered so that $\mathbf{x}_{i}=\mathbf{x}_{i}^{\ast}$ for $i=1,\ldots,m$. We also define the $n\times n$ matrices $U\coloneqq(u_{ij})_{1\leq i,j\leq n}$ and $U^{\prime}\coloneqq(u_{ij}^{\prime})_{1\leq i,j\leq n}$, where $u_{ij}\coloneqq u_{p}(\mathbf{x}_{i},\mathbf{x}_{j})$ and $u_{ij}^{\prime}\coloneqq u_{p}(\mathbf{x}_{i}^{\ast},\mathbf{x}_{j}^{\ast})$. Since $u_{p}$ is a reproducing kernel, $U$ and $U^{\prime}$ are Gram matrices, thus symmetric and positive semi-definite. Furthermore, we define $W_{i}^{0}\coloneqq W_{i}-1$ for any $i=1,\ldots,n$, and define the vectors $\mathbf{V}_{1}\coloneqq(W_{1}^{0},\ldots,W_{m}^{0},0,\ldots,0)^{\top}$ and $\mathbf{V}_{2}\coloneqq(0,\ldots,0,W_{m+1}^{0},\ldots,W_{n}^{0})^{\top}$, so that $\mathbf{W}^{0}=\mathbf{W}-1=\mathbf{V}_{1}+\mathbf{V}_{2}$. It follows that the bootstrap sample $D^{2}_{\mathbf{W}}(\mathbb{X}_{n})$ defined in (3) can be decomposed as
$D^{2}_{\mathbf{W}}(\mathbb{X}_{n})\;$ $=\;\frac{1}{n^{2}}\sum_{1\leq i,j\leq n}W_{i}^{0}W_{j}^{0}u_{ij}\;=\;\frac{1}{n^{2}}(\mathbf{W}^{0})^{\top}U\mathbf{W}^{0}\;=\;\frac{1}{n^{2}}\left(\mathbf{V}_{1}^{\top}U\mathbf{V}_{1}+\mathbf{V}_{2}^{\top}U\mathbf{V}_{2}+2\mathbf{V}_{1}^{\top}U\mathbf{V}_{2}\right)\;.$
@@
where the last line follows from Cauchy-Schwarz inequality applied to the $U$-weighted inner product $(\mathbf{x},\mathbf{x}^{\prime})\mapsto\mathbf{x}^{\top}U\mathbf{x}^{\prime}$ and the $U^{\prime}$-weighted inner product $(\mathbf{x},\mathbf{x}^{\prime})\mapsto\mathbf{x}^{\top}U^{\prime}\mathbf{x}^{\prime}$, which are well-defined since $U,U^{\prime}$ are positive semi-definite. The proof proceeds by bounding each term separately. We discuss how to bound $|\mathbf{V}_{1}U\mathbf{V}_{1}|$ and $|\mathbf{V}_{2}U\mathbf{V}_{2}|$, and the argument for $U^{\prime}$ is identical. We first define the $n\times n$ matrix with $(A_{1})_{ij}=u_{ij}$ for $i,j\leq m$ and $(A_{1})_{ij}=0$, so that we can write $\mathbf{V}_{1}^{\top}U\mathbf{V}_{1}=\mathbf{W}^{\top}A_{1}\mathbf{W}$. We can apply the Hanson-Wright lemma (Lemma˜10) to conclude that there exists positive constants $C$ such that, for any $\delta>0$ and any almost-sure sequence $\mathbb{X}_{n}$, the following event occurs with probability at least $1-\delta/4$,
@@
$|q_{\infty,1-\alpha}(\mathbb{X}_{n})-q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})|\;\leq\;\max(q_{\infty,1-\alpha}(\mathbb{X}_{n}),\;q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast}))\;$ $\leq\;q_{\infty,1-\alpha}(\mathbb{X}_{n})+q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})$
@@ A.3.2 Proof of Lemma˜12
Pick $\{\mathbf{z}_{i}\}_{i=1}^{m^{\prime}}\subset\mathbb{R}^{d}$, and define $\mathbb{Z}_{m^{\prime}}\coloneqq\{\mathbf{z}_{i}\}_{i=1}^{m^{\prime}}$. Decomposing the test statistic $D^{2}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})$ using a similar approach as in (10), we have
@@ A.3.3 Proof of Lemma˜13
Fix any $\theta,\gamma\geq 0$. Let $m^{\prime},n$ be any positive integers with $m^{\prime}\leq f_{n}$, where $f_{n}=o(n^{1/2})$. Define $m=n-m^{\prime}$. Pick $Z_{m^{\prime}}=\{\mathbf{z}_{i}\}_{i=1}^{m^{\prime}}\subset\mathbb{R}^{d}$. The LHS of the inequality can be bounded as
$\;\big{|}\Pr\nolimits_{\mathbb{X}_{m}^{\ast}\sim P,\mathbf{W}}\big{(}\Delta_{\theta}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})>q_{\infty,1-\alpha}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})\big{)}-\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big{(}\Delta_{\theta}(\mathbb{X}_{n}^{\ast})>q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})\big{)}\big{|}$
$=\;\big{|}\Pr\nolimits_{\mathbb{X}_{m}^{\ast}\sim P,\mathbf{W}}\big{(}D(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})>q_{\infty,1-\alpha}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})+\theta\big{)}$
@@
$\big{|}\Pr\nolimits_{\mathbb{X}_{m}^{\ast}\sim P,\mathbf{W}}\big{(}\Delta_{\theta}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})>q_{\infty,1-\alpha}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})\big{)}-\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big{(}\Delta_{\theta}(\mathbb{X}_{n}^{\ast})>q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})\big{)}\big{|}$
$\leq\;\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big{(}\big{|}D(\mathbb{X}_{n}^{\ast})-\gamma_{\ast}\big{|}\;\leq\;t_{m,n,\delta}+\rho_{m,n,\delta}\big{)}+\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big{(}\big{|}D(\mathbb{X}_{n}^{\ast})-\gamma_{\ast}\big{|}\;\leq\;\rho_{m,n,\delta}\big{)}+3\delta$
@@
$\big{|}\Pr\nolimits_{\mathbb{X}_{m}^{\ast}\sim P,\mathbf{W}}\big{(}\Delta_{\theta}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})>q_{\infty,1-\alpha}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})\big{)}-\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big{(}\Delta_{\theta}(\mathbb{X}_{n}^{\ast})>q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})\big{)}\big{|}$
 $\leq\;2\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big{(}\big{|}D(\mathbb{X}_{n}^{\ast})-\gamma_{\ast}\big{|}\;\leq\;\eta(m^{\prime})\big{)}+3\delta$
@@
$\max_{m^{\prime}\leq f_{n}}\sup_{\mathbf{z}_{1},\ldots,\mathbf{z}_{m^{\prime}}\in\mathbb{R}^{d}}\big{|}\Pr\nolimits_{\mathbb{X}_{m}^{\ast}\sim P,\mathbf{W}}\big{(}\Delta_{\theta}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})>q_{\infty,1-\alpha}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})\big{)}$
$\;\qquad\qquad\qquad\qquad\;-\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big{(}\Delta_{\theta}(\mathbb{X}_{n}^{\ast})>q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})\big{)}\big{|}$
$\leq 2\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\left(\big{|}D(\mathbb{X}_{n}^{\ast})-\gamma_{\ast}\big{|}\;\leq\;\eta\left(f_{n}\right)\right)+3\delta$
To complete our proof, it remains to show that the first term can be bounded by $\delta$, from which the claimed result in Lemma˜13 would follow by redefining $\delta$. To proceed, we write

### v5

@@ A.1.3 Proof of Proposition˜9
Proof We first show that the moment conditions (i)-(ii) in Lemma˜7 and Lemma˜8 are met, and apply these lemmas to conclude the proof.
To verify the moment conditions, we first bound the squared Stein kernel evaluated at any $\mathbf{x},\mathbf{z}\in\mathbb{R}^{d}$ as
@@
Having showed the moment conditions, we are now ready to prove the main theorem. Pick any $s\in[0,1)$ and any integer $m^{\prime}\in\big[\frac{1}{2}n^{1-s},n\big)$. Define $\epsilon_{m,n}\coloneqq m^{\prime}/n$ and $m=n-m^{\prime}$. Let $C,C_{P,1},C_{P,2}$ be the constants defined in Lemma˜8, and define the events
@@ A.1.4 Proof of Theorem˜1
Proof It suffices to show the claim for any sequence of the form $\epsilon_{n}=n^{-s}$, where $s\in[0,1)$ is arbitrary. Let the random variable $M^{\prime}$ be defined as at the start of Section˜A.1. For any random sample $\mathbb{X}_{n}$, define the event $\mathcal{A}(\mathbb{X}_{n})=\{D^{2}(\mathbb{X}_{n})>q^{2}_{\infty,1-\alpha}(\mathbb{X}_{n})\}$. For $n$ sufficiently large so that $n^{(1-s)/2}>2\sqrt{2}C\log(1/\alpha)$, we can apply Proposition˜9 to conclude that, for all $n$, there exists $\mathbf{z}_{n}\in\mathbb{R}^{d}$ such that the following holds for all $m^{\prime}\in\big[\frac{1}{2}n^{1-s},n\big)=\big[\frac{1}{2}\epsilon_{n}n,n\big)$,
@@
which will imply the claimed result of our theorem.
Using a concentration inequality for Binomial distributions (e.g., Chung and Lu, 2002, Lemma 2.1) applied to the case of $\text{Binomial}(n,\epsilon_{n})$, the event $\mathcal{B}\coloneqq\{\epsilon_{n}n/2\leq M^{\prime}\leq 3\epsilon_{n}n/2\}$ holds with high probability
@@
$\stackrel{{\scriptstyle(b)}}{{=}}\;\min_{m^{\prime}\in I_{n}}\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}\big(\mathcal{A}(\mathbb{X}_{n})\;|\;M^{\prime}=m^{\prime}\big)\Pr(\mathcal{B})$
@@
where the first inequality follows from the law of total probability, $(a)$ and $(b)$ hold by the partition $\mathcal{B}=\cup_{m^{\prime}\in I_{n}}\{M^{\prime}=m^{\prime}\}$ and again the law of total probability, $(c)$ follows from (24) and (26), and $(d)$ holds since the assumptions $\epsilon_{n}=n^{-s}$ and $s\in[0,1)$ imply $f(n)=o\big(n^{-(1-s)}\big)$. On the other hand, the assumed moment conditions imply $\mathbb{E}_{\mathbf{X}^{\ast}\sim P}[u_{p}(\mathbf{X}^{\ast},\cdot)]=0$ as argued in the paragraph before (11) and $\mathbb{E}_{\mathbf{X}^{\ast}_{1},\mathbf{X}^{\ast}_{2}\sim P}[u_{p}(\mathbf{X}^{\ast}_{1},\mathbf{X}^{\ast}_{2})^{2}]\leq\mathbb{E}_{\mathbf{X}^{\ast}\sim P}[u_{p}(\mathbf{X}^{\ast},\mathbf{X}^{\ast})^{2}]<\infty$ as argued in Section˜A.1.3, so we can apply Arcones and Gine (1992, Theorem 3.5) with $r=2$ to yield the asymptotic validity of the KSD test when $Q=P$, namely
@@ A.1.5 Implications for KSDs on Non-Euclidean Spaces
The core argument underpinning the proofs in Theorem˜1 is that the function $\mathbf{z}\mapsto u_{p}(\mathbf{z},\mathbf{z})$ is unbounded on the support of the model, which in our case is $\mathbb{R}^{d}$. The proof per se does not rely on the support being an Euclidean space. Consequently, this argument can potentially be extended to KSDs constructed for non-Euclidean data, such as Riemannian data (Xu and Matsuda, 2021), sequence data with varying dimensionality (Baum et al., 2023), functional data (Wynne et al., 2022), or censored time-to-event data (Fernandez et al., 2020).
@@ A.2 Proof of Lemma˜2
Proof Direct expansion of the Stein kernel in (1) with a tilted kernel $k$ gives
@@
where the last line follows from Cauchy-Schwarz inequality. The RHS of (29) is bounded over $\mathbf{x}$ assuming the stated conditions on $h$, $w$ and the supremum of $\|s_{p}(x)w(x)\|_{2}=\|s_{p,w}(x)\|_{2}$. To prove the bound on $u_{p}(\mathbf{x},\mathbf{x}^{\prime})$, we denote by $\mathcal{H}_{u}$ the RKHS associated with $u_{p}$, which is a reproducing kernel (see, for example, Barp et al., 2024, Theorem 1). We use the reproducing property of the Stein reproducing kernel $u_{p}$ and the Cauchy-Schwarz inequality for the corresponding RKHS norm $\|\cdot\|_{\mathcal{H}_{u}}$ to yield
@@ A.3 Proof of Theorem˜3 and Related Preliminary Results
  • Section˜A.3.2 shows Lemma˜12, which states that the difference in the exceedance probabilities of $\Delta_{\theta}(\mathbb{X}_{n})$ and of $\Delta_{\theta}(\mathbb{X}_{n}^{\ast})$ is small.
  • Section˜A.3.3 shows Lemma˜13, which states that the rejection probability of a robust-KSD test using $\mathbb{X}_{n}$ is close to that of a test using $\mathbb{X}_{n}^{\ast}$, assuming $\mathbb{X}_{n}$ and $\mathbb{X}_{n}^{\ast}$ differ by no more than $o(n^{1/2})$ elements. Its proof relies on Lemma˜11 and Lemma˜12.
  • Section˜A.3.4 uses Lemma˜13 to show Proposition˜14, which states that the robust-KSD test is qualitatively robust for any $\theta\geq 0$. This result immediately implies Theorem˜3, the proof of which is also contained in this subsection.
@@
Assume $k$ is a tilted kernel satisfying the conditions in Lemma˜2. Then there exists absolute constants $C_{1},C_{2}>0$ such that, for any $\delta>0$ and any (deterministic) realizations $\mathbb{X}_{n}=\{\mathbf{x}_{i}\}_{i=1}^{n}$ and $\mathbb{X}_{n}^{\ast}=\{\mathbf{x}_{i}^{\ast}\}_{i=1}^{n}$ that differ by at most $m^{\prime}$ elements, we have
@@
$\omega(m^{\prime})\;$ $\coloneqq\;\big|\Pr\nolimits_{\mathbb{X}_{m}^{\ast}\sim P,\mathbf{W}}\big(\Delta_{\theta}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})>q_{\infty,1-\alpha}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})\big)$
$\;\quad\quad-\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big(\Delta_{\theta}(\mathbb{X}_{n}^{\ast})>q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})\big)\big|\;.$
@@ A.3.1 Proof of Lemma˜11
Proof Without loss of generality, we assume $\mathbb{X}_{n}$ and $\mathbb{X}_{n}^{\ast}$ are ordered so that $\mathbf{x}_{i}=\mathbf{x}_{i}^{\ast}$ for $i=1,\ldots,m$. We also define the $n\times n$ matrices $U\coloneqq(u_{ij})_{1\leq i,j\leq n}$ and $U^{\prime}\coloneqq(u_{ij}^{\prime})_{1\leq i,j\leq n}$, where $u_{ij}\coloneqq u_{p}(\mathbf{x}_{i},\mathbf{x}_{j})$ and $u_{ij}^{\prime}\coloneqq u_{p}(\mathbf{x}_{i}^{\ast},\mathbf{x}_{j}^{\ast})$. Since $u_{p}$ is a reproducing kernel, $U$ and $U^{\prime}$ are Gram matrices, thus symmetric and positive semi-definite. Furthermore, we define $W_{i}^{0}\coloneqq W_{i}-1$ for any $i=1,\ldots,n$, and define the vectors $\mathbf{V}_{1}\coloneqq(W_{1}^{0},\ldots,W_{m}^{0},0,\ldots,0)^{\top}$ and $\mathbf{V}_{2}\coloneqq(0,\ldots,0,W_{m+1}^{0},\ldots,W_{n}^{0})^{\top}$, so that $\mathbf{W}^{0}=\mathbf{W}-1=\mathbf{V}_{1}+\mathbf{V}_{2}$. It follows that the bootstrap sample $D^{2}_{\mathbf{W}}(\mathbb{X}_{n})$ defined in (3) can be decomposed as
$D^{2}_{\mathbf{W}}(\mathbb{X}_{n})\;=\;\frac{1}{n^{2}}\sum_{1\leq i,j\leq n}W_{i}^{0}W_{j}^{0}u_{ij}\;$ $=\;\frac{1}{n^{2}}(\mathbf{W}^{0})^{\top}U\mathbf{W}^{0}$
$=\;\frac{1}{n^{2}}\left(\mathbf{V}_{1}^{\top}U\mathbf{V}_{1}+\mathbf{V}_{2}^{\top}U\mathbf{V}_{2}+2\mathbf{V}_{1}^{\top}U\mathbf{V}_{2}\right)\;.$
@@
where the last line follows from Cauchy-Schwarz inequality applied to the $U$-weighted inner product $(\mathbf{x},\mathbf{x}^{\prime})\mapsto\mathbf{x}^{\top}U\mathbf{x}^{\prime}$ and the $U^{\prime}$-weighted inner product $(\mathbf{x},\mathbf{x}^{\prime})\mapsto\mathbf{x}^{\top}U^{\prime}\mathbf{x}^{\prime}$, which are well-defined since $U,U^{\prime}$ are positive semi-definite. The proof proceeds by bounding each term separately.
We discuss how to bound $|\mathbf{V}_{1}U\mathbf{V}_{1}|$ and $|\mathbf{V}_{2}U\mathbf{V}_{2}|$, and the argument for $U^{\prime}$ is identical. We first define the $n\times n$ matrix with $(A_{1})_{ij}=u_{ij}$ for $i,j\leq m$ and $(A_{1})_{ij}=0$, so that we can write $\mathbf{V}_{1}^{\top}U\mathbf{V}_{1}=\mathbf{W}^{\top}A_{1}\mathbf{W}$. We can apply the Hanson-Wright lemma (Lemma˜10) to conclude that there exists positive constants $C$ such that, for any $\delta>0$ and any almost-sure sequence $\mathbb{X}_{n}$, the following event occurs with probability at least $1-\delta/4$,
@@
$|q_{\infty,1-\alpha}(\mathbb{X}_{n})-q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})|\;$ $\leq\;\max(q_{\infty,1-\alpha}(\mathbb{X}_{n}),\;q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast}))$
$\leq\;q_{\infty,1-\alpha}(\mathbb{X}_{n})+q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})$
@@ A.3.2 Proof of Lemma˜12
Proof Pick $\{\mathbf{z}_{i}\}_{i=1}^{m^{\prime}}\subset\mathbb{R}^{d}$, and define $\mathbb{Z}_{m^{\prime}}\coloneqq\{\mathbf{z}_{i}\}_{i=1}^{m^{\prime}}$. Decomposing the test statistic $D^{2}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})$ using a similar approach as in (10), we have
@@ A.3.3 Proof of Lemma˜13
Proof Fix any $\theta,\gamma\geq 0$. Let $m^{\prime},n$ be any positive integers with $m^{\prime}\leq f_{n}$, where $f_{n}=o(n^{1/2})$. Define $m=n-m^{\prime}$. Pick $Z_{m^{\prime}}=\{\mathbf{z}_{i}\}_{i=1}^{m^{\prime}}\subset\mathbb{R}^{d}$, and denote the LHS of the inequality by
$T(\mathbf{z}_{1},\ldots,\mathbf{z}_{m}^{\prime})\;$ $\coloneqq\;\big|\Pr\nolimits_{\mathbb{X}_{m}^{\ast}\sim P,\mathbf{W}}\big(\Delta_{\theta}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})>q_{\infty,1-\alpha}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})\big)$
$\;\qquad-\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big(\Delta_{\theta}(\mathbb{X}_{n}^{\ast})>q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})\big)\big|\;.$
The LHS of the inequality can be bounded as
$T(\mathbf{z}_{1},\ldots,\mathbf{z}_{m}^{\prime})\;$ $=\;\big|\Pr\nolimits_{\mathbb{X}_{m}^{\ast}\sim P,\mathbf{W}}\big(D(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})>q_{\infty,1-\alpha}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})+\theta\big)$
@@
$T(\mathbf{z}_{1},\ldots,\mathbf{z}_{m}^{\prime})\;$ $\leq\;\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big(\big|D(\mathbb{X}_{n}^{\ast})-\gamma_{\ast}\big|\;\leq\;t_{m,n,\delta}+\rho_{m,n,\delta}\big)$
$\;\quad+\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big(\big|D(\mathbb{X}_{n}^{\ast})-\gamma_{\ast}\big|\;\leq\;\rho_{m,n,\delta}\big)+3\delta$
@@
$T(\mathbf{z}_{1},\ldots,\mathbf{z}_{m}^{\prime})\;$  $\leq\;2\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big(\big|D(\mathbb{X}_{n}^{\ast})-\gamma_{\ast}\big|\;\leq\;\eta(m^{\prime})\big)+3\delta$
@@
$\max_{m^{\prime}\leq f_{n}}\sup_{\mathbf{z}_{1},\ldots,\mathbf{z}_{m^{\prime}}\in\mathbb{R}^{d}}T(\mathbf{z}_{1},\ldots,\mathbf{z}_{m}^{\prime})\;$ $\leq 2\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\left(\big|D(\mathbb{X}_{n}^{\ast})-\gamma_{\ast}\big|\;\leq\;\eta\left(f_{n}\right)\right)+3\delta$
It remains to show that the first term on the RHS can be bounded by $\delta$, from which the claimed result in Lemma˜13 would follow by redefining $\delta$. To proceed, we write